# Estimating the Number of Street Vendors in New York City††thanks: Thanks to Mohamed Attia, Eric Auerbach, Debipriya

### v1

@@ 4. Discussion
We conclude by discussing the implications of our estimates. In particular, we assess the completeness of the OATH and ACS data. Our estimates suggest both data sets miss the majority of street vendors, and we provide several explanations for this discrepancy.
@@
Ninety-percent of respondents in the Street Vendor Project survey indicated that they live in New York City. Assuming 90% of the estimated 23,000 vendors also live in New York City, we would estimate roughly 20,700 street vendors live in New York City. In comparison, the 2017-2021 ACS estimates 5,997 New Yorkers work the occupation of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers. This suggests that according to the ACS, there are at most 5,997 street vendors in New York City, and thus our finding indicate that the ACS estimate misses 70% or more of street vendors.
@@
A fairer comparison might account for the fact that the standard error of our estimate is approximately 1,000, and the standard error of the ACS estimate is approximately 581, which we obtained from the replicate weights. Using the lower limit of a 95% confidence interval for our estimate (19,000) and the upper 95% confidence limit for the ACS estimate (7,159), our findings suggest the ACS misses roughly two thirds of street vendors. Note the importance of the assumptions stated in Section 3.2 since other derivations of ratio estimation may imply smaller or larger standard errors, from which we might conclude the ACS has lower or higher coverage.
@@
The OATH data indicate that approximately 2,500 (unique) individuals allegedly violate laws related to vending each year. If the ACS estimate of at most 5,997 vendors holds, then nearly half of all vendors are summoned to court each year. If our estimate of 23,000 vendors holds, then approximately one in ten vendors are summoned to court each year. We believe the latter is more realistic.
@@
There may be several reasons why our estimates are significantly larger than the American Community Survey. The fact that the Street Vendor Project distributed relief aid may have encouraged respondents who are unlikely to respond to the ACS. Moreover, the Street Vendor Project, through membership lists, referrals, and canvassing may have identified individuals missed by the Census Bureau’s sampling frame.
@@
Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations. For example, the ACS estimate is the averaged number of vendors between 2017 and 2021, while our estimate reflects 2020 and 2021. For another example, the Census Bureau determines the occupation of an ACS respondent by autocoding responses to write-in questions. It is possible that the Street Vendor Project’s definition of a vendor is more inclusive. Indeed, street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the gig economy.

### v2

@@ 5. Discussion
We conclude by discussing the implications of our estimates. In particular, we assess the completeness of the New York City Office of Administrative Trials and Hearings (OATH) and American Community Survey (ACS) data. Our estimates suggest that while both data sets are spatially representative, they miss the majority of street vendors. We provide several explanations for this discrepancy.
@@
We estimate 23,000 vendors work in New York City. In comparison, the 2017-2021 ACS estimates 4,634 people work the occupation of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in New York City. This suggests that according to the ACS, there are at most 4,634 street vendors in New York City, and thus our finding indicate that the ACS estimate misses 80% of street vendors or more.
@@
A fairer comparison might account for the fact that the standard error of our estimate is approximately 1,000, and the standard error of the ACS estimate is approximately 488, which we obtained from the replicate weights. Using the lower limit of a 95% confidence interval for our estimate (19,000) and the upper 95% confidence limit for the ACS estimate (5,610), our findings suggest the ACS misses roughly two thirds of street vendors or more.
Note the importance of the assumptions stated in Section 3.2 since other derivations of ratio estimation may imply smaller or larger standard errors, from which we might conclude the ACS misses a smaller or larger portion of the street vending population. For example, if the standard errors from Section 4.2 hold, then the confidence intervals would overlap if there were more than 225 vendors in the average market. We believe this is highly unlikely, suggesting that even if the Poisson assumption is violated, the two estimates are inconsistent.
@@
Additional evidence comes from the OATH data. The OATH data indicate that approximately 2,500 (unique) vendors are summoned to court for allegedly violating laws related to vending each year. If the ACS estimate of at most 4,634 vendors holds, then more than half of all vendors are summoned to court each year. If our estimate of 23,000 vendors holds, then approximately one in ten vendors are summoned to court each year. We believe the latter is more realistic.
@@
There may be several reasons why our estimates are significantly larger than the ACS. The fact that the Street Vendor Project distributed relief aid may have encouraged respondents who are unlikely to respond to the ACS. Moreover, the Street Vendor Project, through membership lists, referrals, and canvassing may have covered individuals missed by the Census Bureau’s sampling frame.
@@
Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations. For example, the ACS estimate is the average number of vendors between 2017 and 2021, while our estimate reflects 2020 and 2021. For another example, the Census Bureau determines the occupation of an ACS respondent by autocoding responses to write-in questions. It is possible that the Street Vendor Project’s definition of a vendor is more inclusive. Indeed, street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the gig economy. In this case, either estimate may be relevant, depending on the intended use case.